Raner, Harris & Chan is a consulting firm that specializes in information systems for medical and dental clinics. The firm has two offices—one in Chicago and one in Minneapolis. The firm classifies the direct costs of consulting jobs as variable costs. A contribution format segmented income statement for the company’s most recent year is given:

Office
Total Company Chicago Minneapolis
Sales $ 1,050,000 100.0 % $ 210,000 100 % $ 840,000 100 %
Variable expenses 567,000 54.0 % 63,000 30 % 504,000 60 %
Contribution margin 483,000 46.0 % 147,000 70 % 336,000 40 %
Traceable fixed expenses 235,200 22.4 % 109,200 52 % 126,000 15 %
Office segment margin 247,800 23.6 % $ 37,800 18 % $ 210,000 25 %
Common fixed expenses not traceable to offices 168,000 16.0 %
Net operating income $ 79,800 7.6 %

Required:

1-a. Compute the companywide break-even point in dollar sales.

1-b. Compute the break-even point for the Chicago office and for the Minneapolis office.

1-c. Is the companywide break-even point greater than, less than, or equal to the sum of the Chicago and Minneapolis break-even points?
Answer #1

Answer-1(a):

Dollar sales for company to break even = (Traceable fixed expense + Common fixed expense) / Overall CM ratio

= (235,200 + 168,000) / 0.46 = $876,522

Answer-1(b):

Dollar sales for Chicago office = Segment traceable fixed expense / Overall CM ratio

= 109,200 / 0.70 = $156,000

Dollar sales for Minneapolis office = Segment traceable fixed expense / Overall CM ratio

= 126,000 / 0.40 = $315,000

Answer-1(c):

The sum of the segment break-even points is less than the companywide break-even point because the companywide break-even point takes into account common fixed expenses that do not affect the segment break-even calculations.